OpenStudy (anonymous):

How can you solve the following integrals:

OpenStudy (anonymous):

\[\int\limits_{a}^{b} e ^{2x}/(1+e^{x}) dx\] --> its just an indefinite integral

OpenStudy (anonymous):

\[\int\limits_{0}^{1} xln \left| 1 + x ^{2} \right|\]

OpenStudy (anonymous):

*dx

hartnn (hartnn):

u=1+e^x du = e^x dx e^{2x} = e^x *e^x where e^x = u-1

OpenStudy (anonymous):

yea im stuck at u-1/ u du

hartnn (hartnn):

for 2nd, 1+x^2=u

hartnn (hartnn):

(u-1)/u = u -1/u

OpenStudy (anonymous):

oh so i can integrate u/u - 1/u x.x

OpenStudy (anonymous):

Sorry the x.x isn't part of it

hartnn (hartnn):

yes, you can, easily.

hartnn (hartnn):

for 2nd, you'll just need to integrate ln u

OpenStudy (anonymous):

I thought I needed to use integration by part for expressions that had the product rule :/

hartnn (hartnn):

when you can substitute and simplify, why go for parts ? :)

OpenStudy (anonymous):

oooohh and u= 1+ x^2 and du= 2xdx and du/2 = xdx. that was helpful :D

OpenStudy (anonymous):

so you don't always have to use integration by part when using something that includes the product rule? ( I'm referring to things like xe^x

hartnn (hartnn):

still you'll need integration by parts to integrate ln u

hartnn (hartnn):

not necessary always, forst try substitution

hartnn (hartnn):

*first

OpenStudy (anonymous):

ah so I have to change the limits of integration then go the integration by part. ALright thanks a lot!!

hartnn (hartnn):

thats correct :) welcome ^_^